Grade 11Physical Chemistry

if |z1| = |z2| and arg(z1/z2) = pi then z1+z2 is equal to
a) 0
b) Purely Imaginary
c) Purely Real
d) None
[Ans. a) 0]

Shadic The Hedgehog
6 Years agoGrade 11

2 Answers

Arun
6 Years ago
Let |z1| = |z2| = r ,
and arg(z2) = A
Given, arg(z1) - arg(z2) = π
arg(z1) = π + arg(z2)
arg(z1) = π + A
So, z2 = r cos(A) + i sin(A)
z1 = r cos(π + A) + i sin(π + A)
z1 = - [r cos(A) + i sin (A)]
z1 = - z2
z1 + z2 = 0.
chumtiya
3 Years ago
what the hell!? brody ois sos id hdi fjiw aknfkd iufdkn eda9kjd aojfjis isd jsi sijwi edwmwie isi wi u   iwueuei wiuewiiwe   wdkofjhrbftyoa oz  zxx x xudud  xx x xu ue ejv